### v6

Abstract
The National Statistics Bureau of Ecuador carries out monthly polls to monitor the evolution of the Consumer Price Index, a metric measuring consumer prices of essential commodities. These surveys are administered across a designated set of stores, with a fleet of vehicles transporting pollsters from the bureau headquarters to the chosen locations. Moreover, pollsters move between stores using pedestrian paths or using a vehicle to shorten the travel time. This paper introduces the Integrated Vehicle and Pollster Routing Problem and presents an integer programming model to effectively schedule pollster visits to selected stores while optimizing the routing of the vehicle fleet. Results on the computational complexity, a three-phase algorithm, and computational experience based on real-world instances are provided.
keywords:
routing, scheduling, integer programming, primal heuristics.
1 Introduction
The design of routes for one or several vehicles on a network, satisfying constraints related to the transportation of goods, people, or information, is known as the Vehicle Routing Problem (VRP). On the other hand, crew routing is the process of designing routes for crews to perform specific activities. Real-world applications usually involve vehicle and crew routing tasks that need to be solved in an integrated manner. Considering that crews may be able to interchange vehicles and manage synchronized times, this makes the problem combinatorially complex and more challenging.
This paper focuses on a version of the joint vehicle and crew routing and scheduling problem. The problem arises from data collection procedures designed by the National Statistics Bureau of Ecuador (INEC) to compute the national Consumer Price Index (CPI). A group of pollsters and a fleet of vehicles traverse shared pedestrian and vehicular routes to visit stores scattered throughout the city of Guayaquil and its outskirts over several days. A pollster visits each store to collect the prices of specific goods needed to calculate the CPI. Furthermore, pollsters can walk or be transported by fixed-capacity vehicles during a workday with morning and afternoon shifts, separated by a mandatory break. Every vehicular route must begin at INEC headquarters, dropping off and picking up pollsters and allowing vehicles to return or wait at each store. At the end of each workday, vehicles and pollsters must return to the headquarters. This problem aims to reduce the sum of fixed and variable daily costs. The former is associated with the maintenance of the depot, while the latter is related to hiring pollsters and vehicles.
1.1 Literature review
The Vehicle Routing Problem (VRP) is a well-known and complex combinatorial problem that has received considerable attention since the 1950s. Thus, Dantzig and Ramser [1] formally introduced the VRP as a generalization of the Traveling Salesman Problem, where a simple matching-based heuristic for its solution was proposed. A few years later, Clarke and Wright [2] developed one of the most widely used approaches for this problem with an iterative procedure that enables the rapid selection of an optimal or near-optimal route. Iterative search methods [3, 4, 5] can be easily adapted to problems with a large number of constraints, although they require considerable computing times and several parameter settings. Seminal concepts for exact methods appeared successfully in the 1980s when Christofides et. al. [6] proposed tree search algorithms incorporating lower bounds computed from shortest spanning $k$-degree center trees and $q$-paths. Additionally, Laporte et. al. [7] proposed Gomory cutting planes and branch-and-bound (B&B) approaches based on an integer formulation. Since then, various applications of the VRP have been reported in [8, 9, 10], as well as a variety of exact algorithms [11, 12], heuristics, and metaheuristics [13, 14, 15, 16].
The routing problem addressed in this work arises from two well-established variants of the VRP: the Pickup & Delivery Problem (PDP) and the Dial-a-Ride Problem (DARP). These problems have been studied using a variety of approaches, including exact and heuristic methods.
On the one hand, the PDP constitutes a relevant subset of routing problems in which commodities or passengers have to be transported from different origins to multiple destinations. Savelsbergh and Sol [17] discussed several characteristics distinguishing the PDP from standard VRPs and surveyed these problems and solution methods. Similarly, Parragh et. al. [18] focused on routing problems involving pickups and deliveries related to the transportation of goods from the depot to line-haul customers and from back-haul customers to the depot. Ferrucci [19] considered scenarios where clients can be seen as intermediaries who can also return goods. Sartori and Buriol [20] solved the PDP using a matheuristic based on the Iterated Local Search method with an embedded Set Partitioning Problem that is iteratively solved to recombine routes of previously found solutions. Furtado et. al. [21] proposed a modeling strategy that explicitly assigns vehicles to routes in two-index flow formulations. Wang et. al. [22] developed a variant of the pickup and delivery problem, called the two-echelon pickup and delivery problem using public transport; they formulated a new mathematical model based on a space-time network and adopted an adaptive large neighborhood search (ALNS) algorithm for the problem. Extensive surveys on this problem concerning exact and heuristic approaches are available in [23, 24, 25, 26].
On the other hand, the DARP can be considered a generalization of the PDP and the VRP with Time Windows, as it consists of designing vehicle routes and schedules for individual persons who specify pickup and delivery requests. The most common application can be found in door-to-door transportation services [27, 28]. Furthermore, Madsen et. al. [29] presented a problem characterized by multiple capacities and multiple objectives, describing a system for solving a static Dial-a-Ride routing and scheduling problem with time windows. Braekers et. al. [30] proposed an exact solution approach for small problem instances of the Multi-Depot Heterogeneous DARP. In [31], a dynamic DARP with time window constraints is addressed, along with a two-phase insertion algorithm based on route perturbations. When transportation requests of users between a set of pickup points and a set of delivery points are considered, Masson et. al. [32] provided a solution method based on an ALNS metaheuristic and explained how to check the feasibility of a request insertion. Cortenbach et. al. [33] introduced a mixed-integer linear programming formulation for the Dial-a-Ride Problem with Meeting Points; two preprocessing steps and three valid inequalities are introduced, along with two versions of the Tabu Search metaheuristic, to approximate the optimal solution in large-scale networks. Comprehensive surveys on the DARP can be found in [34, 35, 36].
Joint vehicle and crew routing and scheduling problems are also reported in the literature. For example, Lam et. al. [37] introduced a problem in which crews can interchange vehicles, resulting in space and time interdependencies between vehicle routes and crew routes. They proposed a constraint programming model that overlays crew routing constraints on a standard VRP. In [38], a project motivated by the Austrian Red Cross is discussed. A solution procedure is provided for the daily planning of home healthcare providers operating multiple vehicles to deliver nurses to clients and pick them up after service. Moreover, the authors introduced a matheuristic consisting of two stages: identifying potential walking routes and optimizing the transport system. Molenbruch et. al. [39] introduced a routing algorithm and integrated scheduling procedure to enforce the synchronization of flexible vehicle routes with timetables of public transport services, enabling the design and operational implementation of an integrated mobility system. In the same context, Drexl et. al. [40] studied a simultaneous vehicle and crew routing and scheduling problem arising in long-distance road transport in Europe. A solution heuristic based on a two-stage decomposition of the problem was developed, considering the European Union social legislation for drivers. Lucci et. al. [41] presented a simultaneous vehicle routing and crew scheduling problem (SVRCSP) in which long-distance pickup-and-delivery requests must be fulfilled over a multiday planning horizon, subject to several constraints such as multiple time windows and hour-of-service regulations, among others. The authors developed an algorithm based on the GRASP × ILS metaheuristic. Furthermore, Shen and Li [42] introduced a new model for an integrated electric Vehicle Scheduling Problem (VSP) and Crew Scheduling Problem (CSP). This model employs minimum cost flow formulations for the electric VSP, set partitioning for the CSP, and linking constraints. Due to the nonlinear integer nature of the problem, they proposed a method that combines a matching-based heuristic and an integer linear programming solver.
1.2 Contribution and structure
This paper makes several contributions to the vast literature on Transportation Problems, which is reviewed in the context of this work in the next section. First, a problem, which is a variation of the General Pickup and Delivery Problem (GPDP) and the Dial-a-Ride Problem (DARP), is proposed. Unlike the GPDP, where transshipments at intermediate locations and autonomous load movements are not allowed, this problem considers the possibility that a pollster, left at a specific node to perform a task, might be picked up by another vehicle. Additionally, a pollster can move through the network from one node to another using a pedestrian route. Thus, transshipments and autonomous transportation of the freight are included in this model. This fact highlights the need for syncing vehicular and pedestrian routes while respecting a maximum arrival time corresponding to the length of a working day. Second, this research studies the complexity of the problem, proving it is $\mathcal{NP}$–hard. This result leads to modeling, algorithmic, and computational challenges, which have been addressed using Integer Programming tools and a three-phase heuristic. Finally, the results presented in this study provide a better understanding of pollstering logistics for a public institution and demonstrate a considerable reduction of operational costs of up to 50%.
The structure of this paper is as follows: in Section 2, an IP-based formulation for the Integrated Vehicle and Pollster Routing Problem is introduced, along with a proof of its computational complexity. Due to the hardness of the problem, Section 3 presents a three-phase heuristic. Subsequently, a real-world instance and the computational results are described in Section 4. Finally, concluding remarks are provided in Section 5.
2 An integrated model for routing pollsters and vehicles
2.1 Notation
Let $n\in\mathbb{N}$ be the number of stores, and let $S$ be the set of days in the planning horizon. In this problem, only one depot is considered due to technical requirements. It is assumed that pollsters must depart from and return to the depot in a vehicle on a daily basis. Define $G\coloneqq(V,A)$ as a multi-graph with a node set given by $V\coloneqq\{0,\ldots,2n+1\}$ and an arc set $A\subseteq V\times V$. Each store and the depot are associated with two nodes in $V$, as illustrated in Figure 1. Specifically, the depot is represented by node 00, and its duplicate is node $12n+$ 1. Similarly, the subsets $C_{-}\coloneqq\{1,\dots,n\}$ and $C_{+}\coloneqq\{n+1,\dots,2n\}$ correspond to the stores and their duplicates, respectively. Node duplication enables the modeling of vehicle arrival, service, waiting, and departure times separately. Furthermore, define $C\coloneqq C_{-}\cup C_{+}$, which will collectively be referred to as the set of stores.
Figure 1: Node duplication: The depot, labeled 00, and two stores, labeled $i$ and $j$, are associated with their duplicates, labeled $12n+$ 1, $i+n$, and $j+n$, respectively.
The set $A\coloneqq A_{S}\cup A_{W}\cup A_{V}$ contains three different kinds of arcs. The first group, $A_{S}\coloneqq\big{\{}(i,i+n):i\in C_{-}\big{\}}$, is the set of arcs that connect each store with its duplicate; these are called service arcs. The second group of arcs, $A_{W}\coloneqq\big{\{}(i,j):i\in C_{+},j\in C_{-},i\neq j+n\big{\}}$, called walking arcs, links each pair of stores. Finally, the vehicle transportation arc set, $A_{V}\coloneqq\big{\{}(i,j):i,j\in C,i\neq j,i\neq j+n\big{\}}\cup\big{\{}(0,i):i\in C\big{\}}\cup\big{\{}(i,2n+1):i\in C\big{\}}$, represents potential connections between nodes or depots using vehicles. Moreover, for each arc $a\in A_{S}$, the parameter $t_{a}>0$ represents the time spent by a pollster collecting data at a store; for each arc $a\in A_{W}$, $t_{a}>0$ is the pedestrian travel time between stores; and for each arc $a\in A_{V}$, $\tau_{a}\geq 0$ is the vehicular travel time between two nodes ($\tau_{a}=0$ only if $a\in A_{S}$). A representation of the resulting multigraph is displayed in Figure 2.
Figure 2: Multigraph construction for the IVPRP. Focusing only on the depot and two nodes $i,j\in C_{-}$: there are only outgoing arcs from 00 and only incoming arcs to $12n+$ 1. One service arc is defined for $i$ and $j$ separately, two walking arcs connect $\{i,j,i+n,j+n\}$, and 18 vehicle arcs connect $\{0,i,j,i+n,j+n,2n+1\}$.
Visits are carried out by a set $E$ of pollsters transported by a homogeneous fleet of vehicles $K$ with a positive integer capacity $Q$. Each operational day has an associated fixed cost of $\kappa_{0}\geq 0$. Similarly, vehicles and pollsters have fixed hiring costs of $\kappa_{1}\geq 0$ and $\kappa_{2}\geq 0$, respectively, and they work within a daily time frame $[0,B_{\max}]$. Due to labor regulations, pollsters must take a break of $P\geq 0$ minutes starting within the time interval $[T_{0},T_{1}]\subset[0,B_{\max}-P]$.
A vehicular route is a simple path from node 00 to $12n+$ 1 using exclusively arcs in $A_{V}$. A pedestrian route, also called a walking path, is a simple path from $i\in C_{-}$ to $j\in C_{+}$ consisting of an alternating sequence of arcs in the sets $A_{S}$ and $A_{W}$. A feasible service route for a pollster is defined as a simple path from 00 to $12n+$ 1 linking pedestrian routes with segments of vehicular routes. Thus, at every node $i$, a pollster spends $t_{i,i+n}$ units of time to accomplish a task, while respecting the maximum daily duty length. Finally, a feasible daily plan consists of different feasible service routes synced according to the duration of each pedestrian and vehicular route, and any sequence of feasible daily plans forms an aggregated daily plan.
The Integrated Vehicle and Pollster Routing Problem (IVPRP) consists of finding an aggregated daily plan to visit every node $i\in C$ exactly once at minimum cost. This plan must hire at most $|K|$ vehicles with capacity $Q$ and $|E|$ pollsters within no more than $|S|$ days. The IVPRP is $\mathcal{NP}$–hard, as proved in the following theorem.
Theorem 1.
The IVPRP is $\mathcal{NP}$–hard.
Proof.
The $\mathcal{NP}$–hardness is proved by using a polynomial transformation from the Traveling Salesman Problem (TSP). Let $K_{n+1}$ ($n\geq 3$) be a complete graph with a set of nodes $\{0,1,\dots,n\}$ and weights $c:E(K_{n+1})\to\mathbb{R}_{+}$. An instance of the IVPRP is constructed as follows: an arbitrary node in $V(K_{n+1})$ is chosen as the depot. Without loss of generality, let 00 be such a node. The remaining nodes in $V(K_{n+1})$ correspond to every store in our problem, i.e., $C\coloneqq C_{-}\cup C_{+}$, where $C_{-}=V(K_{n+1})\setminus\{0\}$. The multigraph $D$ is constructed as described above. The cost function $c$ in the TSP corresponds to the vehicular travel times $\tau$ for arcs $(i,j)$, $i,j\in C$, $i\neq j$ in $A_{V}$, and 00 otherwise. Pedestrian travel time is set to $t_{a}\gg\sum_{a\in A_{V}}\tau_{a}$, and $t_{a}=0$ for $a\in A_{S}$. Moreover, fix $|E|=|K|=|S|=1$, $Q=1$, $P=0$, $B_{\max}=\infty$, and the break time window is set to $[0,B_{\max})$.
If the Hamiltonian cycle $\{0,i,j,k,\dots,0\}$ is a feasible solution of the TSP with cost $H$, then the IVPRP has a feasible solution with the same cost. This solution is given by the vehicular route $\{0,i,i+n,j,j+n,k,k+n,\dots,2n+1\}$, the pedestrian route $\bigcup_{i\in C_{-}}(i,i+n)$, and both forming a service route, which is also a feasible daily plan.
Conversely, suppose that the IVPRP has a feasible solution with an aggregated daily plan of length less than $\sum_{a}\tau_{a}$. Since the pedestrian travel times are larger than $\sum_{a}\tau_{a}$, the pedestrian arcs in $A_{W}$ are never used in that solution. Thus, the feasible service route reveals a Hamiltonian cycle for the TSP with the same cost. ∎
The $\mathcal{NP}$–hardness condition for the Integrated Vehicle and Pollster Routing Problem makes it unlikely to find an optimal solution in polynomial time. As a consequence, Section 3 is devoted to identifying heuristic methods to obtain good-quality solutions in a reasonable amount of time.
2.2 Formulation
For the sake of simplicity, the integrated vehicle and pollster routing model (IVPRM) is presented in the following blocks. Table 1 summarizes all the relevant parameters used in the model.
Table 1: Relevant parameters and sets of the proposed model.
Symbol Parameter
$n$ Number of stores
$K$ Set of vehicles
$E$ Set of pollsters
$S$ Set of days in the time horizon
$B_{\max}$ Day length
$P$ Length of the break for pollsters
$Q$ Vehicle capacity
$G\coloneqq(V,A)$ Multi-graph
$V\coloneqq C\cup\{0,2n+1\}$ Node set of $G$
$C\coloneqq C_{-}\cup C_{+}$ Set of nodes associated with stores
$C_{-}\coloneqq\{1,\dots,n\}$ Set of nodes representing the start of service
$C_{+}\coloneqq\{n+1,\dots,2n\}$ Set of nodes representing the end of service
$A\coloneqq A_{S}\cup A_{W}\cup A_{V}$ Arc set of $G$
$A_{S}$ Service arcs connecting each store with its duplicate
$A_{W}$ Walking arcs
$A_{V}$ Vehicle arcs
$t_{a}$ Time spent by a pollster collecting data at a store, for $a\in A_{S}$
$t_{a}$ Pedestrian travel time between stores, for $a\in A_{W}$
$\tau_{a}$ Vehicular travel time between stores, for $a\in A_{V}$
$\kappa_{0}$ Fixed operational cost per working day
$\kappa_{1}$ Fixed vehicle hiring cost
$\kappa_{2}$ Fixed pollster hiring cost
$[T_{0},T_{1}]$ Time window in which pollsters must start their break
$M$ Sufficiently large number
Variables
The following sets of binary variables for arc selection are used in the model. First, variable $x_{i,j}^{e,s}$ is equal to one if and only if the arc $(i,j)\in A_{S}\cup A_{W}$ is chosen to be part of some walking path of pollster $e\in E$ on day $s\in S$. Second, variable $y_{i,j}^{k,s}$ is equal to one whenever arc $(i,j)\in A_{V}$ is part of the route of vehicle $k\in K$ on day $s\in S$. Third, $z_{i,j}^{e,s}$ indicates whether pollster $e\in E$ is transported by some vehicle over arc $(i,j)\in A_{V}$ on day $s\in S$. Fourth, $b_{i}^{e,s}$ (correspondingly $f_{i}^{e,s}$) indicates whether a pollster $e\in E$ starts (finishes) a walking path at node $i\in C_{-}$ ($i\in C_{+}$) on day $s\in S$. Fifth, $w_{i}^{e,s}$ indicates whether a pollster $e\in E$ takes a break at node $i\in C_{-}$ on day $s\in S$. Finally, $u_{s}$ determines whether any workload is assigned to pollsters and vehicles on day $s\in S$ for the final aggregated daily plan. Additionally, for each node $i\in C_{-}$, let $B_{i}$ and $B_{i+n}$ represent the arrival and departure times of a pollster at nodes $i$ and $i+n$, respectively.
Objective function
The objective function aims to minimize the number of operational days and the number of vehicles and pollsters hired. The first component represents the fixed operational daily cost, while the other two components represent the variable costs associated with the number of pollsters and vehicles hired for an operational day.
Although this definition of an objective function is uncommon in Routing Problems, it is justified in this case because the distance walked by a pollster or the travel time spent by a vehicle are irrelevant quantities for this particular application. Instead, the focus is on the fixed daily costs of hiring vehicles and pollsters.
$\min\,\kappa_{0}\,\sum_{s\in S}u^{s}+\kappa_{1}\,\sum_{s\in S}\sum_{k\in K}\sum_{j\in C}y_{0,j}^{k,s}+\kappa_{2}\,\sum_{s\in S}\sum_{e\in E}\sum_{j\in C}z_{0,j}^{e,s}.$ (1)
Pollster routing
The following set of constraints describes the routes for each pollster $e\in E$ within each day $s\in S$ in the time horizon.
$\displaystyle\sum_{s\in S}\sum_{e\in E}x_{i,i+n}^{e,s}$ $=1$ $\forall i\in C_{-},$ (2a)
$\displaystyle\sum_{j\in C_{+}}x_{j,i}^{e,s}-x_{i,i+n}^{e,s}$ $=-b_{i}^{e,s}$ $\forall i\in C_{-},e\in E,s\in S,$ (2b)
$x_{i-n,i}^{e,s}-\displaystyle\sum_{j\in C_{-}}x_{i,j}^{e,s}$ $=f_{i}^{e,s}$ $\forall i\in C_{+},e\in E,s\in S,$ (2c)
$\sum_{j\in C\cup\{0\}}z_{j,i}^{e,s}$ $\leq 1-x_{i,i+n}^{e,s}+b_{i}^{e,s}$ $\forall i\in C_{-},e\in E,s\in S,$ (2d)
$\sum_{j\in C\cup\{2n+1\}}z_{i,j}^{e,s}$ $\leq 1-x_{i-n,i}^{e,s}+f_{i}^{e,s}$ $\forall i\in C_{+},e\in E,s\in S,$ (2e)
$\displaystyle\sum_{j\in C\cup\{0\}}z_{j,i}^{e,s}\quad-\sum_{j\in C\cup\{2n+1\}}z_{i,j}^{e,s}$ $=b_{i}^{e,s}$ $\forall i\in C_{-},e\in E,s\in S,$ (2f)
$\displaystyle\sum_{j\in C\cup\{0\}}z_{j,i}^{e,s}\quad-\sum_{j\in C\cup\{2n+1\}}z_{i,j}^{e,s}$ $=-f_{i}^{e,s}$ $\forall i\in C_{+},e\in E,s\in S,$ (2g)
$\sum_{i\in C}z_{0,i}^{e,s}$ $\leq u_{s}$ $\forall e\in E,s\in S.$ (2h)
Constraints Equation 2a ensure that each store $i\in C_{-}$ is visited exactly once by a pollster on a given day. Constraints Equation 2b and Equation 2c identify the edges of walking paths; specifically, if $b_{i}^{e,s}=1$ (resp. $f_{i}^{e,s}=1$), then a walking path must begin (resp. end) at node $i\in C_{-}$ (resp. $C_{+}$). Moreover, constraints Equation 2dEquation 2g guarantee that if a node $i$ is the beginning or end of a walking path, then exactly one vehicle must pick up or deliver one pollster at that node. Finally, constraints Equation 2h enforce that each pollster departs from the depot at most once on every active day.
Vehicle routing
The following constraints define vehicular routes for pollster transportation.
$\displaystyle\sum_{k\in K}\sum_{j\in C\cup\{2n+1\}}y_{i,j}^{k,s}$ $=\sum_{e\in E}b_{i}^{e,s}$ $\forall i\in C_{-},\,\forall s\in S,$ (3a)
$\displaystyle\sum_{k\in K}\sum_{j\in C\cup\{2n+1\}}y_{i,j}^{k,s}$ $=\sum_{e\in E}f_{i}^{e,s}$ $\forall i\in C_{+},\,\forall s\in S,$ (3b)
$\sum_{i\in C\cup\{0\}}y_{i,j}^{k,s}\quad-\sum_{i\in C\cup\{2n+1\}}y_{j,i}^{k,s}$ $=0$ $\forall j\in C,k\in K,s\in S,$ (3c)
$\displaystyle\sum_{j\in C}y_{0,j}^{k,s}-\sum_{j\in C}y_{j,2n+1}^{k,s}$ $=0$ $\forall k\in K,s\in S,$ (3d)
$\sum_{i\in C}y_{0,i}^{k,s}$ $\leq u_{s}$ $\forall k\in K,s\in S,$ (3e)
$\sum_{e\in E}z_{i,j}^{e,s}$ $\leq Q\sum_{k\in K}y_{i,j}^{k,s}$ $\forall(i,j)\in A,s\in S.$ (3f)
Constraints Equation 3a and Equation 3b ensure that vehicles arrive only at pickup nodes and depart only from delivery nodes. Constraints Equation 3c and Equation 3d are flow conservation constraints. Furthermore, Equation 3e ensure that each vehicle departs from the depot at most once on every active day. Finally, Equation 3f establish the linkage between pollster and vehicular routes.
Time management and shift length
The following constraints are introduced to define time bounds for the duration of a pollster shift as well as the times required to travel and visit each store.
$B_{i+n}\geq B_{i}+t_{i,i+n}+P\sum_{s\in S}\sum_{e\in E}w_{i}^{e,s}$ $\forall i\in C_{-},$ (4a)
$B_{j}\geq B_{i}+t_{i,j}-M\bigg{(}1-\sum_{s\in S}\sum_{e\in E}x_{i,j}^{e,s}\bigg{)}$ $\forall i\in C_{+},j\in C_{-},$ (4b)
$B_{j}\geq B_{i}+\tau_{i,j}-M\bigg{(}1-\sum_{s\in S}\sum_{k\in K}y_{i,j}^{k,s}\bigg{)}$ $\forall(i,j)\in A_{V},i\neq 0,j\neq 2n+1,$ (4c)
$B_{i}\geq\tau_{0,i}-M\bigg{(}1-\sum_{s\in S}\sum_{k\in K}y_{0,i}^{k,s}\bigg{)}$ $\forall i\in C,$ (4d)
$B_{\max}u_{s}\geq B_{i}+\tau_{i,2n+1}-M\left(1-\sum_{k\in K}y_{i,2n+1}^{k,s}\right)$ $\forall i\in C,s\in S,$ (4e)
Constraints Equation 4a require that the departure time from a store $i$ be at least as large as the arrival time plus the service time and, if applicable, the time needed for a pause at that store. Note that the sum in the right-most term of these constraints equals one if a pause is taken at node $i$, and zero otherwise. Constraints Equation 4b and Equation 4c ensure that the arrival time at a store within a route accounts for the pedestrian or vehicular travel time from the previously visited store, respectively. If node $i$ immediately follows the depot in any route, then constraints Equation 4d provide a lower bound on the arrival time of vehicles and pollsters. Finally, constraints Equation 4e link the variables $B_{i}$ and $u^{s}$ to bound the latest arrival time of any vehicle at the depot for each day $s$. Here, $M$ is a sufficiently large number such that $M>\max\{t,\tau\}+B_{\max}$.
Pollster breaks
The following constraints ensure that each pollster on duty takes a break starting within the prescribed time frame on a given day.
$T_{0}\sum_{s\in S}\sum_{e\in E}w_{i}^{e,s}\leq B_{i}+t_{i,i+n}\leq T_{1}+M\bigg{(}1-\sum_{s\in S}\sum_{e\in E}w_{i}^{e,s}\bigg{)}$ $\forall i\in C_{-},$ (5a)
$w_{i}^{e,s}\leq x_{i,i+n}^{e,s}$ $\forall i\in C_{-},e\in E,s\in S,$ (5b)
$\sum_{i\in C_{-}}w_{i}^{e,s}=\sum_{j\in C}z_{0,j}^{e,s}$ $\forall e\in E,s\in S.$ (5c)
Constraints Equation 5a establish that if a break is taken by a pollster at node $i$, it must occur within the time interval $[T_{0},T_{1}]$. Constraints (5b) guarantee that a pollster takes a break only at visited stores, and constraints Equation 5c enforce that each pollster on duty takes a break exactly once per day.
2.3 Symmetry breaking inequalities
In any feasible aggregated daily plan, a fixed vehicle services a vehicular route; however, the route can be served by any other vehicle in $K$, yielding the same objective value. Similarly, the workload of any pair of pollsters in $E$ can be permuted within a daily plan. Moreover, the same applies to the set of days $S$ concerning the feasible daily plans. As a result, a set of equivalent solutions might arise by relabeling assigned vehicles and pollsters. The following set of constraints aims to avoid these symmetries. This approach can be seen as an application of the well-known technique of symmetry elimination [43, 44, 45].
$\sum_{i\in C}z_{0,i}^{0,0}=1,$ (6a)
$\sum_{i\in C}z_{0,i}^{e,s}\leq\sum_{i\in C}z_{0,i}^{e-1,s},$ $\forall e\in E\setminus\{0\},s\in S,$ (6b)
$\sum_{i\in C}z_{0,i}^{e,s}\leq\sum_{i\in C}z_{0,i}^{e,s-1},$ $\forall e\in E,s\in S\setminus\{0\},$ (6c)
$\sum_{i\in C}y_{0,i}^{0,0}=1,$ (6d)
$\sum_{i\in C}y_{0,i}^{k,s}\leq\sum_{i\in C}y_{0,i}^{k-1,s},$ $\forall k\in K\setminus\{0\},s\in S,$ (6e)
$\sum_{i\in C}y_{0,i}^{k,s}\leq\sum_{i\in C}y_{0,i}^{k,s-1},$ $\forall k\in K,s\in S\setminus\{0\},$ (6f)
$\sum_{e\in E}b_{i}^{e,s}\leq 1-\sum_{e\in E}b_{i}^{e,s-1},$ $\forall i\in C_{-},s\in S\setminus\{0\},$ (6g)
$\sum_{e\in E}f_{i}^{e,s}\leq 1-\sum_{e\in E}f_{i}^{e,s-1},$ $\forall i\in C_{+},s\in S\setminus\{0\},$ (6h)
$\displaystyle\sum_{e\in E}\sum_{j\in C_{+}}x_{j,i}^{e,s}\leq 1-\sum_{e\in E}b_{i}^{e,s-1},$ $\forall i\in C_{-},s\in S\setminus\{0\},$ (6i)
$\displaystyle\sum_{e\in E}\sum_{j\in C_{-}}x_{i,j}^{e,s}\leq 1-\sum_{e\in E}f_{i}^{e,s-1},$ $\forall i\in C_{+},s\in S\setminus\{0\},$ (6j)
$\sum_{i\in C_{-}}b_{i}^{e,s}\leq|C_{-}|-\sum_{r=0}^{s-1}\sum_{i\in C_{-}}b_{i}^{e-1,r},$ $\forall e\in E\setminus\{0\},s\in S\setminus\{0\},$ (6k)
$\sum_{i\in C_{+}}f_{i}^{e,s}\leq|C_{+}|-\sum_{r=0}^{s-1}\sum_{i\in C_{+}}f_{i}^{e-1,r},$ $\forall e\in E\setminus\{0\},s\in S\setminus\{0\},$ (6l)
$u^{s}\leq u^{s-1},$ $\forall s\in S\setminus\{0\}.$ (6m)
Constraints Equation 6aEquation 6c ensure an ordered assignment of pollsters, starting with the one labeled as 00, for each day $s\in S$. Similarly, constraints Equation 6dEquation 6f guarantee an ordered assignment for vehicles.
Constraints Equation 6gEquation 6j define daily relationships between served stores across two consecutive days; i.e., if node $i$ is visited on day $s-1$, then the related variables for this node on day $s$ are bounded to zero. Constraints Equation 6kEquation 6l impose an upper bound on the number of stores that were not visited in the previous days. Finally, constraints Equation 6m reinforce the activation of consecutive active days.
2.4 Upper bounds
An upper bound for the objective function of IVPRP can be obtained by solving the following nonlinear problem:
$\min\kappa_{0}\ell_{S}+\kappa_{1}\ell_{K}\ell_{S}+\kappa_{2}\ell_{E}\ell_{S}$ (7a)
subject to
$\dfrac{1}{\ell_{E}}\Big{(}P\ell_{E}\ell_{S}+\sum_{i\in C_{-}}t_{i,i+n}\Big{)}+\dfrac{1}{\ell_{K}}\sum_{i\in V}\Big{[}\max_{(i,j)\in A_{V}}\tau_{i,j}+\max_{(j,i)\in A_{V}}\tau_{j,i}\Big{]}\leq B_{\max}\ell_{S},$ (7b)
$\ell_{K}\leq\ell_{E}\leq Q\ell_{K},$ (7c)
$1\leq\ell_{S}\leq|S|,\qquad 1\leq\ell_{K}\leq|K|,\qquad 1\leq\ell_{E}\leq|E|,$ (7d)
$\ell_{S},\ell_{K},\ell_{E}\in\mathbb{N};$ (7e)
where $\ell_{S}$, $\ell_{K}$, and $\ell_{E}$ are decision variables representing the number of needed days, vehicles, and pollsters, respectively. The optimal solution $(\ell_{S}^{*},\ell_{K}^{*},\ell_{E}^{*})$ of Equation 7aEquation 7d provides an upper bound on the objective function Equation 1 of model IVPRM. As a result, this solution determines a tighter value for the cardinality of the sets $S$, $K$, and $E$ required to obtain an aggregated daily plan.
Observe that constraint Equation 7b divides a day into two portions. On the one hand, the first term averages the pollstering time and breaks that each pollster would take across all days to complete an aggregated daily plan. On the other hand, the second term averages the time required, in a worst-case scenario, for vehicles to travel, drop off pollsters, and pick them up at each node. The sum of both terms in any feasible solution for IVPRM must necessarily be below the time horizon $B_{\max}$. Furthermore, constraint Equation 7c limits the number of pollsters according to vehicle capacity. Finally, the optimal solution can be obtained, in the worst case, using exhaustive enumeration in $\mathcal{O}\big{(}|S||K||E|\big{)}$ time.
2.5 Lower bounds
Likewise, it is possible to determine lower bounds based on store visiting times. To do so, first, let $\ell_{e}$ be the smallest integer value representing the minimum number of pollsters required, in at most $|S|$ days, to distribute the aggregated service time while considering the maximum shift length and mandatory break durations. This value is obtained by solving the following nonlinear program:
$\max\bigg{\{}\frac{1}{\ell_{e}}\sum_{i\in C_{-}}t_{i,i+n}:\,\frac{1}{\ell_{e}}\sum_{i\in C_{-}}t_{i,i+n}\leq B_{\max}-P\wedge\ell_{e}\in\{1,\ldots,|S||E|\}\bigg{\}}.$ (8)
Notice that if problem (8) is infeasible, then there are not enough resources to solve IVPRM, thereby providing an infeasibility certificate.
Next, the minimum number of days required to visit all stores, denoted as $\ell_{s}$, is recovered as the first index $s$ such that $\ell_{e}$ belongs to the set $\Lambda_{E,s}=\{s,s+1,\ldots,s\cdot|E|\}$ for all $s\in\{1,\ldots,|S|\}$. Here, $\Lambda_{E,s}$ is the set containing the cumulative number of pollsters that can be active on a day with index $s$.
Finally, a bound on the number of vehicles is derived by noting that if a day is active, then there must also be an active vehicle (see Equation 3f), and each vehicle can transport at most $Q$ pollsters. Thus, the smallest number $\ell_{k}$ that satisfies these constraints solves the following linear program: $\min\big{\{}\ell_{k}:\,\ell_{e}\leq Q\ell_{k}\,\wedge\,\ell_{k}\in\{\ell_{s},\ldots,\ell_{s}|K|\}\big{\}}$.
A solution for these three problems can be obtained through exhaustive enumeration in $\mathcal{O}\big{(}|S||E|+|K|\big{)}$ time. Moreover, while the upper bounds $(\ell_{S},\ell_{K},\ell_{E})$ would change depending on the costs in the objective function, the lower bounds $(\ell_{s},\ell_{k},\ell_{e})$ are cost-independent. These lower bounds are incorporated into IVPRM as the following constraints:
$\sum_{s\in S}u^{s}$ $\geq\ell_{s},$ (9a)
$\sum_{s\in S}\sum_{k\in K}\sum_{j\in C}y_{0,j}^{k,s}$ $\geq\ell_{k},$ (9b)
$\sum_{s\in S}\sum_{e\in E}\sum_{j\in C}z_{0,j}^{e,s}$ $\geq\ell_{e}.$ (9c)
Combining these inequalities yields the lower bound $\kappa_{0}\ell_{s}+\kappa_{1}\ell_{k}+\kappa_{2}\ell_{e}$ with respect to the objective of IVPRM.
2.6 An illustrated example
To explain the solution approach for the IVPRP, a small instance is presented in Figure 3 with the following parameters: $n=4$ stores, $|E|=2$ pollsters, $|K|=2$ vehicles with a capacity of $Q=2$, and a limit of $|S|=2$ days. The break is fixed at one minute within a time frame of $[T_{0},T_{1}]=[20,30]$, and the time horizon is set to 30 minutes. The fixed and variable costs are $\kappa_{0}=\$300$, $\kappa_{1}=\$100$, and $\kappa_{2}=\$80$, respectively.
Figure 3: Multigraph and optimal solution for an instance of IVPRP.
The vehicular and pedestrian travel times are shown in Table 2. The diagonal of the pollster travel time matrix corresponds to the service times. Note that this table does not include information related to nodes in the set $\{5,6,7,8,9\}$, which are the duplicated nodes explained in the multigraph construction.
Table 2: Vehicle and walking travelling times within each pair of stores.
Vehicle ($\tau$) Pollster ($t$)
00 $1$ $2$ $3$ $4$ $1$ $2$ $3$ $4$
Store 00 $1$ $4$ $6$ $2$
$1$ $2$ $1$ $2$ $10$ $1$ $2$ $3$ $14$
$2$ $1$ $1$ $4$ $15$ $2$ $17$ $8$ $20$
$3$ $7$ $2$ $5$ $9$ $3$ $8$ $1$ $15$
$4$ $1$ $15$ $11$ $9$ $15$ $20$ $15$ $1$
As shown in Figure 3, the optimal solution for this small instance consists of one vehicular route: $0\to 1\to 4\to 8\to 3\to 7\to 6\to 9$ (continuous lines), and three pedestrian routes: $1\to 5\to 2\to 6$ for pollster 00, and $4\to 8$ and $3\to 7$ for pollster $1$. Synced pedestrian and vehicular routes form feasible service routes corresponding to the optimal daily plan. Breaks are represented with the dashed nodes $2$ and $3$ for pollsters 00 and $1$, respectively. The optimal value is achieved in one day using one vehicle and two pollsters, with a total cost of \$$\$560$.
3 Solution approach
3.1 Three-phase algorithm
Given that the IVPRP is $\mathcal{NP}$–hard, preliminary tests have shown that finding an optimal solution for instances with even a small number of nodes cannot be guaranteed in a reasonable amount of time. Therefore, a heuristic procedure is proposed in this subsection. The method consists of three phases solved consecutively.
In the first phase, the nodes are partitioned into an even number of balanced subsets. Each partition is associated with a part-time daily plan, either in the morning or afternoon, without breaks. The second phase takes the graph induced by the nodes of each partition as input and solves the IVPRM problem. This input-reduced version of the main problem is denoted by (R-IVPRM). The final phase adequately links the solutions found for each partition to construct feasible daily plans.
Other approaches using graph-partitioning techniques have been similarly reported in the literature; see [20, 46].
Partitioning
To divide the node set into smaller subsets, the following problem is solved: let $G=(C_{-},F)$ be an undirected complete graph with edge costs corresponding to the pedestrian travel times between each pair of stores. Additionally, let $t:C_{-}\longrightarrow\mathbb{R}^{+}$ be a weight function representing the service times at stores, and let $k\geq 2$ be a fixed integer. A $k$-partition of $G$ is a collection of $k$ subgraphs $(C_{1},F(C_{1})),\ldots,(C_{k},F(C_{k}))$ of $G$, where $\{C_{1},\ldots,C_{k}\}$ forms a partition of $C_{-}$. Moreover, let $W_{L},W_{U}\in\mathbb{R}^{+}$, with $W_{L}\leq W_{U}$, be the lower and upper bounds for the total service time of each subgraph, defined as the sum of the service times of all nodes within the subgraph. The problem consists of finding a $k$-partition such that $W_{L}\leq\sum_{j\in C_{i}}t_{j}\leq W_{U}$ for all $i\in\{1,\ldots,k\}$, while minimizing the total edge cost over all subgraphs.
To solve this partitioning problem, the approach proposed in [47] is used, where the size constraints are relaxed. Hence, $W_{L}$ and $W_{U}$ are computed as $\nicefrac{{\mu n}}{{k}}-\sigma$ and $\frac{\mu n}{k}+\sigma$, respectively, where $\mu$ and $\sigma$ are the sample mean and standard deviation of the service times associated with all the stores.
Routing pollsters and vehicles
For any partition obtained in the previous phase, a reduced version of IVPRM is solved to find feasible vehicular and pedestrian routes. The reduced version, denoted as R-IVPRM, consists of the original model with $|S|=1$, $P=0$, and the removal of constraints Equation 5aEquation 5c. Additionally, simplified versions of the symmetry-breaking inequalities, which do not consider consecutive days, are included; specifically, Equation 6aEquation 6b, and Equation 6dEquation 6e.
A solution of the relaxed LP of IVPRM is used to compute lower bounds on the optimal objective value, and this value should be as tight as possible. To improve these lower bounds, some valid inequalities are derived below.
Remark 1.
The following expressions are valid inequalities for R-IVPRM:
$\frac{1}{|E|}\sum_{i\in C_{-}}(B_{i+n}-B_{i})$ $\leq B_{\max},$ (10a)
$B_{i}$ $\leq B_{i+n},$ $\forall i\in C_{-}.$ (10b)
Inequality Equation 10a is valid because the latest arrival time to the depot is greater than the minimum service time required to collect all data using the maximum number of pollsters; i.e., $\sum_{i\in C_{-}}{t_{i,i+n}}\big{/}{|E|}\leq B_{\max}$. On the other hand, the total time spent at each store is greater than or equal to the total time required to collect data; i.e., $\sum_{i\in C_{-}}t_{i,i+n}\leq\sum_{i\in C_{-}}(B_{i+n}-B_{i})$.
Although constraints Equation 4bEquation 4e imply Equation 10b, they are not necessarily valid when subcycles appear in fractional solutions of the linear relaxation. Hence, as in a feasible solution, it is only possible to visit node $i+n$ after visiting $i$, and $12n+$ 1 must be the last visited node. Therefore, Equation 10b holds.
Linking partitions
Since each partition corresponds to a part-time daily plan, the purpose of this phase is to match two of these subsets into a full-time feasible daily plan. Note that, as each part-time daily plan may involve a different number of active vehicles and pollsters, it can be assumed that any vehicles or pollsters not needed for the afternoon shift can remain at the depot. Conversely, if the workforce is insufficient to fulfill the shift, additional vehicles or pollsters may join the workforce. To avoid inconsistencies, breaks are assigned at the depot between each half-shift, as each day can be split into two half-shifts by setting a new time frame upper bound $\tilde{B}_{\max}=(B_{\max}-P)/2$.
Thus, the following two cases arise for consideration: (a) the required number of vehicles or pollsters differs between at least two sets, and (b) the required number of vehicles and pollsters is the same for all sets in the partition. For the former, the cost-differing sets are linked by matching their costs in descending order by pairs. For the latter, let $\mathcal{K}=(V,E)$ be a complete graph, where each node in $V=\{1,\dots,k\}$ is associated with a feasible solution of R-IVPRM for any subset obtained during the Partitioning procedure. A cost for each edge $\{i,j\}\in E$ is defined as the sum of the arrival times at the depot for the last vehicle in partitions $i$ and $j$. Then, a minimum-cost matching in $\mathcal{K}$ is found. Note that if ties arise in case (a), they are resolved as a special instance of case (b). In either case, a feasible aggregated daily plan is returned.
Feasible daily plan construction
Feasible daily plans are constructed according to Algorithm 1, which connects the phases explained previously.
Input: an instance of IVPRM, maximum solver time $T_{\max}>0$, maximum number of iterations of the partitioning algorithm $N_{\max}\in\mathbb{N}$, an even number $\underline{k}\in\mathbb{N}$.
Output: a feasible solution of IVPRM.
Initialization: $k=\underline{k}$; $f=\texttt{false}$.
while $f=\texttt{false}$ do
     Perform $N_{\max}$ iterations of the partitioning procedure to find a $k$-way partition
     of $V(G)$.
     for $i\in\{1,\ldots,k\}$ do
         Find the best feasible solution of R-IVPRM induced by subset $i$, within the
         time limit $T_{\max}$.
         if R-IVPRM induced by subset $i$ is infeasible then
              $f=\texttt{false}$
              break
         else
              $f=\texttt{true}$
         end if
     end for
     if $f$ is false and $k+2\leq 2|S|$ then
         $k=k+2$
     else
         break
     end if
end while
if $f$ is true then
     Perform the linking procedure.
end if
Algorithm 1 Constructing feasible aggregated daily plans
4 Computational experiments
In this section, computational experiments are presented. These consist of a set of random instances that test Algorithm 1 against attempting to solve IVPRM exactly. Additionally, the resolution of a real-world instance is addressed at the end of the section.
The MIP formulations of IVPRM and R-IVPRM have been implemented in Gurobi 12 [ARM] using the Python programming language. The tests were performed on a MacBook Pro with an M1 processor and 16GB of memory. All instances, as well as the source code of the implementation, can be downloaded from:
4.1 Simulated instances
Random instances were generated using data from the OpenStreetMaps API with the osmnx package [48]. Several Points of Interest (e.g., entertainment venues and shops) were sampled from an urban neighborhood, covering a $1000$ m radius from its center. For each pair of stores, the shortest vehicular and pedestrian paths between them were computed. The average travel time between stores was then calculated using an average speed of $30$ km/h for vehicles and $5$ km/h for pollsters. Since vehicles must respect the street network directions, the resulting vehicle travel time matrix is asymmetric. Store-polling times, in minutes, were also randomly generated from a uniform distribution $\mathcal{U}(1,20)$.
As presented in Table 3, 10 instances were simulated with different values of $n$, $B_{\max}$ in minutes, $P$ in minutes, and $Q$. The number of available vehicles, pollsters, and days to solve each instance was computed using the nonlinear model Equation 7aEquation 7e. This was done by initializing $|K|=3$, $|E|=5$, and $|S|=15$, corresponding to the available real-world resources. Finally, the costs for the objective function were set to $\kappa_{0}=\$200$, $\kappa_{1}=\$100$, and $\kappa_{2}=\$40$.
Table 3: Parameters associated with simulated instances.
Parameter Symbol Values
##\## Stores $n$ 10 12 14 16 18 20 25 30 40 50
##\## Vehicles $|K|$ 3 2 2 2 2 3 2 2 2 3
##\## Pollsters $|E|$ 3 2 2 3 3 4 2 2 4 5
##\## Days $|S|$ 1 2 2 2 2 1 2 2 2 2
Day length $B_{\max}$ 100 100 100 100 100 150 150 200 200 250
Pause $P$ 10 10 10 10 10 15 15 20 20 25
Capacity $Q$ 1 1 1 2 2 2 3 3 4 4
Here, $|K|$ was bounded by $3$, $|E|$ by $5$, and $|S|$ by $15$.
Four different strategies are used to solve the IVPRP, each focusing on a distinct approach to the problem. The first strategy involves solving the Mixed-Integer Linear Program (MIP) IVPRM using Gurobi in its default configuration. This B&B-based method runs for up to 3 00030003\,0003 000 seconds and aims to find optimal solutions when possible. The second procedure combines the B&B approach with Gurobi’s NoRel heuristic. The NoRel heuristic is a general-purpose technique for MIPs that skips solving the LP relaxation and instead uses methods like rounding and presolve reductions to identify feasible solutions. It is run for $120$ seconds, attempting to generate initial solutions. The third approach is the problem-tailored heuristic described in Algorithm 1. This approach is specifically designed for the structure of IVPRM and is particularly effective for large-scale instances due to its partitioning step. The fourth and final strategy is the Fixing Variables Heuristic. This method combines elements of the exact approach and the tailored heuristic. It starts by solving a relaxed version of IVPRM to find the best feasible or optimal solution for an initial partition. The integer variables from this solution are then fixed as inputs for IVPRM, except for those corresponding to the ending nodes $\{0,2n+1\}$. With the fixed variables, the model is solved again within a limited runtime to search for an integer feasible solution. If no solution is found, the process continues with the next partition, iteratively fixing and adding variables until the best feasible or optimal solution is achieved. This strategy is most suitable for smaller instances as it involves solving IVPRM, which does not scale well. The results of these strategies are summarized in Tables 4 and 5.
Table 4: Results of applying MIP model & Gurobi NoRel Heuristic
$n$ Bounds IVPRM NoRel Heuristic
$(\ell_{k},\ell_{e},\ell_{s})$ Low Up $(\hat{k},\hat{e},\hat{s})$ Obj. GAP $(\hat{k},\hat{e},\hat{s})$ Obj. GAP
10 (2,2,1) 480 620 (2,2,1) 480 0.0% (2,2,1) 480 0.0%
12 (2,2,1) 480 960 (2,2,1) 480 0.0% (2,2,1) 480 0.0%
14 (2,2,1) 480 960 (3,3,2) 820 41.5% (3,3,2) 820 41.5%
16 (1,2,1) 380 1040 (2,4,2) 760 50.0%
18 (1,2,1) 380 1040 (3,4,2) 860 55.8%
20 (1,2,1) 380 660 (2,4,1) 560 32.1%
25 (1,2,1) 380 960 (2,4,2) 760 50.0%
30 (1,2,1) 380 960 (2,4,2) 760 50.0%
40 (1,3,1) 420 1120 (4,8,2) 1120 62.5%
50 (2,6,2) 840 1400
Each row represents the results for an instance of size $n$ described in Table 3.
Table 4 compares the behavior of the Gurobi solver under different configurations as mentioned above. The first column presents the number of nodes in the instances. The following three columns, grouped in the multicolumn Bounds, report the bounds obtained following the procedures described in Sections 2.4 and 2.5. The other two multicolumns, IVPRM and NoRel Heuristic, each consist of three columns displaying the triplet $(\hat{k},\hat{e},\hat{s})$, the objective value, and the gap of the solution. Here, the triplet $(\hat{k},\hat{e},\hat{s})$ represents the cummulative number of vehicles used, the number of active pollsters, and the number of days in the plan, respectively. These quantities are computed as follows:
$\hat{s}\coloneqq\sum_{s\in S}u^{s},\qquad\hat{k}\coloneqq\sum_{s\in S}\sum_{k\in K}\sum_{j\in C}y_{0,j}^{k,s},\qquad\hat{e}\coloneqq\sum_{s\in S}\sum_{e\in E}\sum_{j\in C}z_{0,j}^{e,s}.$
The gap is computed using the formula
$\mathrm{GAP}=\frac{|\mathtt{Objective}-\mathtt{Low}|}{\mathtt{Objective}},$
where $\mathtt{Objective}$ is the value of Equation 1, and $\mathtt{Low}$ is the maximum between the lower bound obtained from Section 2.5 and the lower bound obtained from the branch-and-bound process.
Table 5 evaluates the behavior of the two heuristic methods, namely, Algorithm Algorithm 1 and the Fixing Variables Heuristic. The instance information in this table coincides with that of Table 4.
Table 5: Results of applying different solution strategies on the simulated instances.
$n$ Bounds Algorithm 1 Fixing Variables
Heuristic
$(\ell_{k},\ell_{e},\ell_{s})$ Low Up $(\hat{k},\hat{e},\hat{s})$ Obj. GAP $(\hat{k},\hat{e},\hat{s})$ Obj. GAP
10 (2,2,1) 480 620 (2,2,1) 480 0.0% (2,2,1) 480 0.0%
12 (2,2,1) 480 960 (2,2,1) 480 0.0% (2,2,1) 480 0.0%
14 (2,2,1) 480 960 (4,4,2) 960 50.0% (3,3,2) 820 41.5%
16 (1,2,1) 380 1040 (2,4,2) 760 50.0% (2,3,1) 520 26.9%
18 (1,2,1) 380 1040 (2,4,2) 760 50.0% (2,4,2) 760 50.0%
20 (1,2,1) 380 660 (2,3,1) 520 26.9% (2,3,1) 520 26.9%
25 (1,2,1) 380 960 (2,3,2) 720 47.2% (2,3,2) 720 47.2%
30 (1,2,1) 380 960 (2,3,2) 720 47.2% (2,3,2) 720 47.2%
40 (1,3,1) 420 1120 (2,5,2) 800 47.5% (2,5,2) 800 47.5%
50 (2,6,2) 840 1400 (3,7,2) 980 14.3% (3,7,2) 980 14.3%
Each row represents the results for an instance of size $n$ described in Table 3.
Interesting insights can be drawn from the results. First, it is noteworthy that the customized invocation of the NoRel heuristic contributed significantly to finding feasible solutions in all but one instance. However, when comparing the NoRel heuristic with Algorithm 1 and the Fixing Variables Heuristic, NoRel outperforms Algorithm 1 only in instance 14. For all other instances, the proposed algorithms achieve better solutions than NoRel. It is important to remark that the Fixing Variables Heuristic produces a marginal improvement when compared to Algorithm 1.
For completeness, a detailed report on the computational results of Algorithm 1 is provided in Table 6. Each set of stores is partitioned into $2$ or $4$ subsets, labeled A through D. For each subset, the table includes the number of vehicles used $\bar{k}$, the number of active pollsters $\bar{e}$, the cost of the half shift, and the optimality gap. The first two quantities are computed as
$\bar{k}\coloneqq\sum_{k\in K}\sum_{j\in C}y_{0,j}^{k}\qquad\text{and}\qquad\bar{e}\coloneqq\sum_{e\in E}\sum_{j\in C}z_{0,j}^{e}.$
Only partitions where R-IVPRM had a feasible solution for all subsets are included. A final column is provided with the cumulative solution times (in seconds).
Table 6: Results of R-IVPRM for simulated instances at each subset.
$n$ Subset A Subset B Subset C Subset D Time
$(\bar{k},\bar{e})$ Obj. GAP $(\bar{k},\bar{e})$ Obj. GAP $(\bar{k},\bar{e})$ Obj. GAP $(\bar{k},\bar{e})$ Obj. GAP
10 (2,2) 280 0.0% (2,2) 280 0.0% 13.02
12 (2,2) 280 0.0% (2,2) 280 0.0% 1.80
14 (2,2) 280 0.0% (2,2) 280 0.0% (2,2) 280 0.0% (2,2) 280 0.0% 3.16
16 (1,2) 180 0.0% (1,2) 180 0.0% (1,2) 180 0.0% (1,2) 180 0.0% 302.81
18 (1,2) 180 0.0% (1,2) 180 0.0% (1,2) 180 0.0% (1,2) 180 0.0% 111.07
20 (2,2) 280 0.0% (2,3) 320 43.8% 415.54
25 (1,2) 180 0.0% (1,1) 140 0.0% (1,1) 140 0.0% (1,2) 180 22.2% 579.07
30 (1,1) 140 0.0% (1,1) 140 0.0% (1,2) 180 22.2% (1,1) 140 0.0% 448.34
40 (1,2) 180 0.0% (1,3) 220 18.2% (1,2) 180 0.0% (1,2) 180 0.0% 709.54
50 (1,3) 220 18.2% (2,3) 320 43.8% (1,4) 260 30.8% (1,3) 220 18.2% 2 160.97
Each row represents the results for an instance of size $n$ described in Table 3.
4.2 A real-world case study
The case study proposed by INEC focuses on the city of Guayaquil. As the country’s main commercial port, its relevance in economic and population development makes this city of 3 million inhabitants ideal for collecting data to monitor the CPI. Hence, INEC deploys a group of at most $5$ pollsters and $3$ vehicles monthly to collect data from $820$ stores in Guayaquil and its outskirts (see Figure 4). Currently, the planning is done empirically using all available resources and data collection is completed in $17$ days.
The objective of the operational managers at INEC was to reduce the costs associated with the number of vehicles and pollsters within a time horizon of at most $15$ days. Furthermore, each day, the workforce must complete its duties within $7$ hours, including a $40$-minute break. Accordingly, the baseline parameters and sets are $n=820$, $S=\{0,\ldots,14\}$, $E=\{0,1,2,3,4\}$, $K=\{0,1,2\}$, $B_{\max}=420$ minutes, and $P=40$ minutes. Additionally, the cost parameters are set to $\kappa_{0}=\$200$, $\kappa_{1}=\$100$, and $\kappa_{2}=\$40$.
The travel time between each pair of stores was computed using the shortest path in the road and walking networks provided by a geographic information system (GIS). In the walking network, the average walking velocity of pollsters is fixed at $5$ km/h, whereas the average vehicle displacement velocity in the city is fixed at $30$ km/h. In contrast to the simulated instances, a velocity of $40$ km/h is fixed for the outskirts of the city. Service times for each store were measured and provided by INEC.
Computation of the lower bounds, as described in Section 2.5, for the resulting instance of IVPRM yields that any feasible aggregated daily plan requires at least $3$ days, $15$ pollsters, and $4$ vehicles.
Figure 4: Map of Guayaquil showing the locations of stores to be visited.
4.3 Results and solution visualization
Upon the first attempt to solve the IP implementation of IVPRM for a small number of nodes in the graph ($n=15$), no feasible solution was found, even when using an HPC server. Similarly, the variable fixing heuristic also failed to produce a solution. This outcome is not surprising, as the full model contains nearly 400 million variables, resulting in an LP that exceeds available memory. These challenges motivated the development and application of the three-phase solution approach described Algorithm 1.
The following tables summarize the behavior of this algorithm using real-world data. Notably, a lower bound for the number of partitions is determined as the ratio between the total service time and the total working time available per pollster each day ($\underline{k}=10$).
Table 7: Statistics of the partitions obtained with Algorithm 1.
##\## Subsets $\mathbf{n}_{\textbf{min}}$ $\mathbf{n}_{\textbf{max}}$ Minimum Maximum ##\## Subsets with
$k$ service time service time feasible solutions
10 56 100 538.7 570.6 0
12 36 97 446.5 478.1 0
14 16 79 380.7 412.1 0
16 35 65 330.9 362.6 0
18 14 68 292.8 324.0 0
20 11 60 261.6 293.2 1
22 10 57 236.2 268.1 3
24 11 54 215.2 247.0 7
26 10 52 197.5 229.1 6
28 7 43 134.9 214.1 16
30 7 43 153.7 216.4 30
[…]
Even though the maximum number of pollsters and vehicles available at INEC are $|E|=5$ and $|K|=3$, Algorithm 1 consistently reported that two pollsters and one vehicle are sufficient to construct every half-shift of a feasible daily plan. Consequently, Table 8 presents the time of the last vehicular arrival at the depot for each subset in the reported partition with $k=30$. For instance, the workload of subset $24$ is completed in at most $2$ hours. It was observed that half of the instances were solved to optimality in less than two minutes, while the remaining instances required the maximum allowed time.
Table 8: Arrival times for each subset of stores within a $30$–way partition.
Subsets 0 — 9 Subsets 10 — 19 Subsets 20 — 29
Arrival time Arrival time Arrival time
0 1 h 36 min 10 2 h 28 min 20 3 h 03 min
1 1 h 55 min 11 2 h 14 min 21 2 h 01 min
2 2 h 02 min 12 1 h 39 min 22 1 h 43 min
3 1 h 55 min 13 2 h 19 min 23 1 h 45 min
4 2 h 08 min 14 1 h 55 min 24 2 h 00 min
5 1 h 57 min 15 3 h 30 min 25 2 h 08 min
6 2 h 18 min 16 2 h 42 min 26 2 h 18 min
7 1 h 50 min 17 2 h 14 min 27 2 h 18 min
8 2 h 49 min 18 2 h 16 min 28 1 h 47 min
9 3 h 03 min 19 2 h 04 min 29 1 h 54 min
The results of the linking partition step are presented in Table 9. A pair of subsets is assigned to each day $s\in\{0,\ldots,14\}$, corresponding to a morning and afternoon half-shift required to accomplish the workload. For example, day $11$ syncs subsets $27$ and $1$ for a total daily working time of $4$ hours and $53$ minutes, including a $40$-minute break between shifts. The first subset of stores is polled in the morning, and the second is visited in the afternoon.
As shown in Table 8, each daily plan is executed by $2$ pollsters and $1$ vehicle, completing the workload within a time horizon of $15$ days. This result contrasts with the empirical solution provided by INEC, which required $5$ pollsters and $3$ vehicles within a time horizon of $17$ days. Given that the daily income of each pollster is \$$\$40$ and the daily cost of hiring and using a vehicle is \$$\$100$, the empirical solution currently used by INEC approximately doubles the cost of the aggregated daily plan found with Algorithm 1.
Table 9: Daily plan results of Algorithm 1.
Daily plans 0 — 7 Daily plans 8 — 14
$s$ Synced Working $s$ Synced Working
subsets time subsets time
0 15 $\to$ 0 5 h 45 min 8 20 $\to$ 12 5 h 22 min
1 9 $\to$ 22 5 h 26 min 9 8 $\to$ 23 5 h 14 min
2 16 $\to$ 28 5 h 10 min 10 10 $\to$ 7 4 h 59 min
3 13 $\to$ 29 4 h 53 min 11 27 $\to$ 1 4 h 53 min
4 6 $\to$ 14 4 h 54 min 12 26 $\to$ 3 4 h 53 min
5 18 $\to$ 5 4 h 53 min 13 11 $\to$ 24 4 h 55 min
6 17 $\to$ 21 4 h 56 min 14 4 $\to$ 2 4 h 51 min
7 25 $\to$ 19 4 h 52 min Average: 5 h 03 min
Finally, the resulting real aggregated daily plan is depicted in Figure 5. In the upper left corner, the afternoon shift of the eighth working day (subset 12) is zoomed in. The stores visited by the first and second pollsters are represented by circles and squares, respectively, with each store labeled by a number between $1$ and $820$. Continuous and dashed lines indicate the routes: continuous lines represent pedestrian routes, while dashed lines represent vehicle routes.
Figure 5: Map of Guayaquil displaying the partition subsets, stores, and routes.
5 Conclusions
The Integrated Vehicle and Pollster Routing Problem (IVPRP), focused on the transportation of pollsters with pedestrian and vehicular routes, has been introduced in this paper. It generalizes the Pickup & Delivery Problem, combining well-known problems such as the Share-a-Ride, the Integrated Dial-a-Ride, and the periodic Vehicle Routing Problem with time windows and backhaulings. The complexity of the IVPRP has been shown to be $\mathcal{NP}$–hard by reduction from the Traveling Salesman Problem.
A mixed integer formulation was proposed for this problem and assessed using simulated instances and real-world data. Due to the hardness of the problem, a three-phase heuristic method was devised: the set of stores to be visited is partitioned at the first stage, then pollsters and vehicles are routed on each subset, and finally, pairs of subsets are linked as half-shifts of an aggregated daily plan.
The algorithm was tested with a real-world instance at Guayaquil, Ecuador’s second most important city. INEC provided the data with information related to 820 stores to be visited monthly. The heuristically-found solution improved the one designed by INEC, approximately reducing twice the operational costs since half of the available workforce and fleet is required. Moreover, the resulting aggregated daily plan takes two days less than the former empirical solution.
Finally, future research could concentrate on the development of more advanced heuristic and metaheuristic approaches to address larger instances of the IVPRP. This could involve hybrid methods that integrate exact and heuristic techniques to achieve a balance between solution quality and computational efficiency. Additionally, there is significant potential to explore extensions of the IVPRP incorporating more complex constraints, such as dynamic routing adjustments based on real-time data or multi-period planning that accounts for demand variability over time. While this study focuses on a specific case in Ecuador, future research could investigate how the proposed framework might be adapted to similar applications in other fields, including logistics, urban planning, and disaster response.
[…]
Acknowledgments The authors are grateful with the managers of INEC, specially with Alexandra Enríquez for providing the data used to build the real-world instance, and for her valuable feedback on the results. Additionally, we acknowledge the collaboration of Pablo Zuleta during the first stage of this research. We greatly appreciate the anonymous referees’ suggestions, which improved significantly the clarity and impact of the findings of this paper.
[…]
Funding This project was funded by Escuela Politécnica Nacional Research Project PIJ–15-12 with support of the Laboratory of Scientific Computing of the Research Center on Mathematical Modelling – ModeMat (hpcmodemat.epn.edu.ec). Andrés Miniguano Trujillo acknowledges support of MAC-MIGS CDT Scholarship under EPSRC grant EP/S023291/1.
[…]
Conflict of interest/Competing interests The authors have no conflicts of interest to declare that are relevant to the content of this article.